# Ab Initio Theory of Large Phonon Magnetic Moments Induced by Electron-Phonon Coupling in Magnetic Materials

### v3

@@ Abstract
Chiral phonons, characterized by nonzero angular momenta and magnetic moments, have attracted extensive attention. However, a long-standing critical issue in this field is the lack of an approach to accurately calculate phonon magnetic moments resulting from electron-phonon coupling (EPC) in realistic materials. Here, based on the linear response framework, we develop an ab initio theory for calculating EPC-induced magnetic properties of phonons, applicable to both insulating and metallic materials. Then, we evidently demonstrate EPC-induced large phonon magnetic moments and phonon Zeeman splittings resulting from spin-phonon coupling in magnetic metals. Interestingly, these splittings can open substantial topologically nontrivial phonon gaps, generating intrinsic phonon Chern states. Further, by constructing an inertially decoupled lattice model, we predict candidate materials exhibiting such intrinsic phonon Chern states with robust edge phonon currents which are proposed to detect neutral particles, such as dark matter particles. Our work enables ab initio calculations for EPC-induced magnetic properties of phonons and long-sought magnetic phonon spectra.
@@
Introduction. Chiral phonons, collective vibrations stemming from the circular motion of atoms in solids with nonzero angular momentum, were initially proposed by Zhang and Niu [1, 2] and subsequently verified experimentally in monolayer WSe2 [3]. Recently, there has been a surge of research interest in chiral phonons, covering both theoretical investigations [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18] and experiments [19, 20, 21, 22, 23, 24, 25, 26, 27]. Of particular importance are phonon magnetic moments (PMMs) and phonon Zeeman splittings, which play a crucial role in chiral-phonon-involved effects, such as ultrafast Einstein-de Haas Effect [20] and Barnett effect [22, 28]. However, classical models based on point-charge approximations yield PMM values that are orders of magnitude below experimental measurements [11, 19]. While electron-phonon coupling (EPC) has been indicated as a promising route to enhance PMMs [8, 11, 15], an ab initio approach capable of accurately calculating EPC-induced PMMs in realistic materials is lacking.
@@
Figure 1: Schematic of the phonon Zeeman splitting driven by (a) the classical PMM and (b) the EPC-enhanced PMM. In (a), the phonon band crossing point consists of two degenerate phonon modes vibrating in two different directions (black arrows), which can be recombined into left-circularly (red) and right-circularly (blue) polarized chiral phonons. The classical phonon Zeeman splitting deduced from the point-charge model ($Z_{\text{eff}}$ is the effective charge) is negligible as a result of the significant atomic mass, whereas the EPC-induced phonon Zeeman splitting (b) is orders of magnitude larger. This significant splitting can open a substantial topologically nontrivial phonon gap, giving rise to the phonon Chern state with topological edge states [dashed lines in (b)].
Besides magnetic moments, phonons can also exhibit Hall effects [29, 30, 31, 32] and topological characteristics, such as Berry curvature and Chern numbers [33, 34, 35]. When the time-reversal symmetry (TRS) is broken, phonon states with nonzero Chern numbers host a phonon analog of quantum anomalous Hall states, featuring topologically protected phonon edge currents. Nevertheless, TRS-breaking external fields for phonons, such as strong magnetic fields or Coriolis forces, often face experimental limitations [36, 37]. Consequently, phonon Chern states driven by the phonon Zeeman splitting resulting from spin-phonon coupling (SPC) in magnets without requiring external fields are highly desirable. However, it is challenging to predict such intrinsic topological phonon states in magnets due to the inability to perform ab initio calculations of TRS-breaking phonon spectra [38, 39, 40, 41, 42, 43].
@@
To resolve these challenges, we establish an ab initio theory for determining EPC-induced magnetic properties of phonons, which enables first-principles calculations of previously inaccessible TRS-breaking phonon spectra across material classes from insulators to metals. Based on this theory, we demonstrate that EPC-induced PMMs can lead to large phonon Zeeman splittings through SPC in magnetic metals, which are orders of magnitude stronger than classical predictions from point-charge models [Fig. 1(a)]. Considering that these splittings could open significant topological gaps in phonon dispersions [Fig. 1(b)], we propose an inertially decoupled lattice model to realize intrinsic phonon Chern states in realistic materials. Furthermore, through first-principles calculations, we predict a family of promising candidate materials exhibiting intrinsic phonon Chern states. Our work establishes an ab initio theoretical foundation for investigating EPC-induced magnetic properties of phonons, which paves the way for further developments of phonon physics.
@@
Ab initio theory for phonon Zeeman energy and phonon magnetic moment. Previous theories regarding magnetic properties of phonons driven by EPC have mainly been applicable to insulators or semimetals, and are restricted to the $\Gamma$ point [8, 11, 13, 15]. Since EPC is typically much more significant in metallic systems, we develop an ab initio theory that can generally calculate EPC-induced magnetic properties of phonons across the whole Brillouin zone for both metals and insulators, which is also compatible with previous theories [see the supplemental material (SM) [44]]. We will demonstrate our theory as follows.
@@
Instead of directly calculating the magnetic moment of a single phonon, we focus on experimentally observable phonon Zeeman energy induced by the coupling between PMM and spins of electrons. We start from the total phonon Hamiltonian including the TRS-breaking perturbation as
@@
where $a,b=x,y,z$ represent the Cartesian coordinates, $l=1,2,...N$ denotes the lattice site for a system containing $N$ atoms, and $u_{l,a}=\sqrt{m}x_{l,a}$ represents the renormalized atomic displacement with $m$ being the atomic mass. The last term in Eq. (1), denoted as $\delta H$, describes the TRS-breaking perturbation resulting from SPC [1], and $\eta_{l,ab}$ is a $3\times 3$ real antisymmetric matrix satisfying $\eta_{l,ab}=-\eta_{l,ba}$. In addition, $p_{l,a}=\dot{u}_{l,a}+\sum_{l,b}\eta_{l,ab}u_{l,b}$ represents the canonical momentum [44]. By identifying the phonon angular momentum as $s^{\text{ph}}_{l,c}=\sum_{ab}\epsilon_{abc}u_{l,a}p_{l,b}$ (c is the direction of angular momentum), with $\epsilon_{abc}$ being the third-order Levi-Civita tensor, we can express $\delta H$ as phonon Zeeman energy in a compact form $E_{\textrm{Z}}=\bm{\eta}\cdot\bm{s}^{\text{ph}}=\sum_{l,ab}\eta_{l,c}s^{\text{ph}}_{l,c}$, where we have defined $\eta_{l,c}=\frac{1}{2}\sum_{ab}\epsilon_{abc}\eta_{l,ab}$. Notably, we find that the angular momentum of chiral phonon can be treated as an effective perturbative field to the electron-phonon coupled system, and thus the Zeeman energy of chiral phonon can be effectively regarded as a linear response.
Importantly, we will show that $\bm{\eta}$ can be derived from the TRS-breaking self-energy correction of chiral phonons. Within the framework of linear response theory, $\bm{\eta}$ is concisely expressed as $\bm{\eta}(\bm{q},\omega)=\frac{1}{i\omega}\bm{\chi}^{-}(\bm{q},\omega)$, where $\chi_{c}^{-}(\bm{q},\omega)=\chi_{c,\uparrow}(\bm{q},\omega)-\chi_{c,\downarrow}(\bm{q},\omega)$ [see SM [44]], indicating that the spin polarization of electrons leads to Zeeman splittings of chiral phonons. The susceptibility $\chi_{c,s}(\bm{q},\omega)$ for $s=\uparrow,\downarrow$ can be obtained through the Feynman diagram illustrated in Fig. 1(b),
@@
Here, $\bm{k}$ and $\bm{q}$ represent the wave vectors of the electron and phonon, respectively, and $\omega$ is the phonon frequency. $N_{\bm{k}}$ is the total number of $\bm{k}$, and $\epsilon_{s}^{\bm{k},\alpha}$ denotes the electronic band structure with band indices $\alpha$. $f$ is the Fermi-Dirac distribution, and $V_{c,s}^{\bm{k},\alpha\beta}(\bm{q},\omega)$ is the EPC tensor [45, 46]. It should be emphasized that all parameters in Eq. (2) can be directly obtained from first-principles calculations, without empirical inputs or additional assumptions. A non-zero $\bm{\chi}^{-}$ requires spin-polarized electronic states, for example, in ferromagnets. To realize large $\bm{\chi}^{-}$, there are two key points according to Eq. (2): (i) The contribution of EPC tensor $V_{c,s}^{\bm{k},\alpha\beta}(\bm{q},\omega)$ is substantial, and materials composed of light atoms tend to possess strong $V_{c,s}^{\bm{k},\alpha\beta}(\bm{q},\omega)$. (ii) When the energy difference $(\epsilon^{\bm{k}+\bm{q},\alpha}-\epsilon^{\bm{k},\beta})$ approaches $\hbar\omega$, it leads to an enhancement of $\bm{\chi}^{-}$. Consequently, we can conclude that EPC-induced large phonon Zeeman splittings are most likely to occur in ferromagnetic narrow-gap semiconductors, semimetals and metals containing light elements.
@@
Here, $\mu_{B}$ denotes the Bohr magneton of electron, $B_{c}$ is the effective magnetic field, and $g^{\text{ph}}_{c}=\frac{\hbar}{2i\omega}\frac{\partial\chi_{c}^{+}}{\partial E_{F}}g^{e}_{c}$ represents the effective Landé $g$-factor of the chiral phonon (see SM [44]), where $E_{F}$ is the Fermi energy and $\chi_{c}^{+}(\bm{q},\omega)=\chi_{c,\uparrow}(\bm{q},\omega)+\chi_{c,\downarrow}(\bm{q},\omega)$. $g^{e}_{c}$ is the $g$-factor along $c$ direction of electrons involved in EPC. Equation (3) enables first-principles calculations of PMM. The dimensionless parameter $\frac{\hbar}{2i\omega}\frac{\partial\chi_{c}^{+}}{\partial E_{F}}$ acts as an amplification factor dominated by the EPC tensor $V_{c,s}^{\bm{k},\alpha\beta}$ and density of electron states near the Fermi energy [Eq. (S22) in SM [44]], and it allows the $g$-factor of the chiral phonon to reach the order of $g^{e}_{c}$, leading to a large PMM. This enhancement of PMM originates from electron loop currents induced by the EPC-mediated angular momentum transfer from ions to electrons, where electron loop currents are significantly enhanced owing to the tiny electron mass. In contrast, within the classical framework, the phonon’s $g$-factor is constrained to $10^{-5}\sim 10^{-4}$ order of $g^{e}_{c}$ due to large ionic masses compared to the electron mass [11, 47].
Figure 2: Inertially decoupled model (IDM) and intrinsic phonon Chern states. (a) Schematic of chiral phonon modes in the high frequency regime: When $m\sim M$, both atoms contribute to high-frequency modes (upper panel); in the IDM condition ($m\ll M$), high-frequency modes are primarily contributed by light atoms (lower panel). (b) Phonon dispersions of nonmagnetic honeycomb lattice and ferromagnetic triangular lattice. (c-e) TRS-breaking phonon dispersions of the 2D IDM by combining the honeycomb and triangular lattices (c,d) and the 3D IDM by stacking 2D IDM along the $z$-direction (e). (f) Edge states (cyan/magenta) of the 2D IDM. Parameters are detailed in the SM [44].
@@
Intrinsic phonon Chern states and the inertially decoupled model. The accurate calculation of phonon Zeeman splitting enables the determination of magnetic phonon spectra throughout the Brillouin zone, which is indispensable for exploring magnetic phonon topology. Following the analogy to the Chern state in the Haldane model [48], we adopt the two-dimensional (2D) honeycomb lattice [49, 50, 51] as an illustrative example. When either inversion symmetry ($I$) or TRS ($T$) is broken, each phonon Dirac cone at $K$($K^{\prime}$) valley acquires a finite gap, effectively gaining masses denoted as $\Delta_{I}$ and $\Delta_{T}$, respectively. It has been shown that when $|\Delta_{T}|>|\Delta_{I}|$, band inversion occurs in one of the two Dirac cones, resulting in a topological phonon state with a Chern number of $|C|=1$ [49].
@@
However, it is challenging to realize such phonon Chern states in realistic materials with honeycomb lattices. Although Coriolis forces and strong magnetic fields can produce a nonzero $\Delta_{T}$ [49], these approaches either require large external fields or complex experimental measurements. In contrast, the phonon Zeeman splitting could offer a sizable $\Delta_{T}$ in magnets without the need of external fields. To further satisfy the condition of $|\Delta_{T}|>|\Delta_{I}|$, it is necessary to minimize $|\Delta_{I}|$. Therefore, an elemental ferromagnetic honeycomb material inherently preserving inversion symmetry with $|\Delta_{I}|=0$ would be the ideal candidate. Unfortunately, to our knowledge, no elemental ferromagnetic honeycomb material has been discovered. Furthermore, any honeycomb lattice composed of multiple elements could introduce a substantial $\Delta_{I}$, which will make the condition $|\Delta_{T}|>|\Delta_{I}|$ difficult to satisfy.
@@
To seek realistic materials that exhibit intrinsic phonon Chern states, we propose a practical composite lattice model, referred to as the inertially decoupled model that comprises two types of atoms with a significant mass (inertia) difference. A prototypical example of this model is constructed by combining a honeycomb lattice composed of nonmagnetic light atoms of mass $m$ with a triangular lattice of heavy magnetic atoms of mass $M$ ($m\ll M$), as shown in Figs. 2(b,c). Due to the difference in atomic inertia, the high-frequency and low-frequency dispersions are mainly contributed by the light and heavy atoms, respectively, as shown in Figs. 2(a,c), which is termed “inertially decoupled” (see SM [44] for details). Within the honeycomb sublattice, the inversion symmetry is preserved, resulting in a vanishing $\Delta_{I}$ mass at both the $K$ and $K^{\prime}$ valleys. Meanwhile, the magnetic $M$ atoms break TRS to generate large Zeeman splittings of chiral phonons via SPC, producing the desired large $\Delta_{T}$ mass. The large $\Delta_{T}$ opens a topologically nontrivial phonon gap (seen in Fig. 2(d)), thereby realizing an intrinsic phonon Chern state with $|C|=1$, characterized by topologically protected edge states crossing the phonon gap, as shown in Fig. 2(f).
@@
Figure 3: Lattice structure and TRS-breaking phonon dispersions of ferromagnetic $\mathrm{EuSi_{2}}$. (a,b) The front view (a) and the side view (b) of the lattice structure of $\mathrm{EuSi_{2}}$. (c,d) The phonon dispersions of bulk $\mathrm{EuSi_{2}}$ without (c) and with (d) the SPC, respectively. (e) The phonon dispersion of bulk $\mathrm{EuSi_{2}}$ under open-boundary conditions. (f,g) The phonon dispersion of the sandwiched slab enclosed by the black rectangle in (b) without (f) and with (g) the SPC, respectively. (h) The phonon dispersion of the sandwiched slab under open-boundary conditions. The color gradient from red to blue represents the phonon angular momentum. Cyan and magenta colors in (c) and (h) label topological boundary modes residing on opposite boundaries.
Candidate materials. Building upon our theory, we developed a dedicated computational module within Quantum ESPRESSO [52, 53] incorporating density-functional perturbation theory with Hubbard corrections (DFPT+U) [45, 46] for precise determination of response coefficients $\bm{\eta}$ and Landé $g$-factors of phonons. Then we add the effect of $\bm{\eta}$ into Schrödinger-like phonon equations constructed from the dynamical matrix through near-degenerate perturbation theory (see SM [44] for details), yielding TRS-breaking phonon dispersions.

### v4

# Ab Initio Theory of Phonon Magnetic Moment Induced by Electron-Phonon Coupling in Magnetic Materials

@@ Abstract
Circularly polarized phonons, characterized by nonzero angular momenta and magnetic moments, have attracted extensive attention. However, a long-standing critical issue in this field is the lack of an approach to accurately calculate phonon magnetic moments resulting from electron-phonon coupling (EPC) in realistic materials. Here, based on the linear response framework, we develop an ab initio theory for calculating EPC-induced magnetic properties of phonons, applicable to both insulating and metallic materials. Our method can precisely calculate phonon Zeeman splittings in magnetic metals with significant EPC, as demonstrated by the remarkable agreement with recent experimental observations of phonon Zeeman splitting in the ferromagnetic Weyl semimetal $\mathrm{Co_{3}Sn_{2}S_{2}}$. In addition, the long-sought magnetic phonon spectra across the entire Brillouin zone are obtained, facilitating the study of magnetic phonon transport and topology. Specifically, by constructing an inertially decoupled lattice model, we propose candidate materials exhibiting intrinsic phonon Chern states with robust unidirectional edge phonon currents. Our work paves the way for investigating novel phonon phenomena in magnetic quantum materials.
@@
Introduction. Circularly polarized phonons, collective vibrations stemming from the circular motion of atoms in solids that carry nonzero angular momentum [1, 2], are referred to as axial phonons [3]. Recently, there has been a surge of research interest in circularly polarized phonons [4, 5], covering both theoretical investigations [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27] and experiments [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39]. Of particular importance is the phonon magnetic moment (PMM), which is derived from phonon energy shift in a magnetic field [14]. PMM plays a crucial role in various effects involving circularly polarized phonons, such as ultrafast Einstein-de Haas Effect [30] and Barnett effect [32, 40]. However, classical models based on point-charge approximations yield PMM values that are orders of magnitude below experimental measurements [16, 29]. While electron-phonon coupling (EPC) has been indicated as a promising route to enhance PMMs [14, 16], to our knowledge, there is no comprehensive ab initio framework to quantitatively predict EPC-induced PMMs in realistic materials. Previous theories describing EPC-induced PMM are only applicable to either insulators or semimetals [14, 15, 16, 13, 19, 21, 22, 20], and more importantly, most of them rely on empirical parameters and, consequently, deviate from ab initio approach.
@@
Figure 1: Schematic of the phonon Zeeman splitting driven by (a) the classical PMM and (b) the EPC-enhanced PMM. In (a), the phonon band crossing point consists of two degenerate phonon modes vibrating in two different directions (black arrows), which can be recombined into left-circularly (red) and right-circularly (blue) polarized phonons. The classical phonon Zeeman splitting deduced from the point-charge model ($Z_{\text{eff}}$ is the effective charge) is negligible as a result of the significant atomic mass, whereas the EPC-induced phonon Zeeman splitting (b) is orders of magnitude larger.
Meanwhile, the coupling between PMM and magnetic order through spin-phonon interactions yields rich phenomena, such as phonon Zeeman splittings, phonon Hall effects [41, 42, 43, 44], and magnetic topological phonon states [45, 5]. For example, phonon Zeeman splitting induced by magnetic order was recently observed experimentally in the ferromagnetic Weyl semimetal $\mathrm{Co_{3}Sn_{2}S_{2}}$ [46], yet a theory capable of quantitatively characterizing such Zeeman splittings is lacking. Moreover, investigations of phonon transport and phonon topology in magnetic systems usually require the full notion of the magnetic phonon spectra across the entire Brillouin zone. A prominent example is the phonon Chern state (phonon analogue of quantum anomalous Hall effect), which necessitates integrating Berry curvatures throughout the Brillouin zone to correctly obtain the topological Chern number [47, 48, 49, 50, 51]. Although nonmagnetic phonon topology has been comprehensively investigated [52, 53, 54, 55, 56, 57], magnetic phonon topology remains largely unexplored due to challenges of calculating magnetic phonon spectra.
@@
To address these challenges, we establish an ab initio theory for determining EPC-induced magnetic properties of phonons, which enables ab initio calculations of previously inaccessible magnetic phonon spectra across material classes from insulators to metals. Based on this theory, we demonstrate that EPC-induced PMMs can lead to large phonon Zeeman splittings in magnetic metals [Fig. 1(b)], which are orders of magnitude stronger than classical predictions from point-charge models [Fig. 1(a)]. Notably, our theory is further validated by the remarkable agreement with the experimental result of the phonon Zeeman splitting of $\mathrm{Co_{3}Sn_{2}S_{2}}$. By applying our theory, we predict a family of promising candidate materials with phonon Chern states according to the proposed inertially decoupled lattice model. Our work establishes an ab initio theoretical foundation for investigating EPC-induced magnetic properties of phonons, which paves the way for further developments of phonon physics.
@@
Ab initio theory for phonon Zeeman energy and phonon magnetic moment. Previous theories addressing magnetic properties of phonons driven by EPC have mainly been applicable to insulators or semimetals, and are restricted to the $\Gamma$ point [14, 16, 13, 19]. Since EPC is typically much more significant in metallic systems, we develop an ab initio theory that can generally calculate EPC-induced magnetic properties of phonons across the whole Brillouin zone for both metals and insulators, which is also compatible with previous theories [see the supplemental material (SM) [58]]. We will demonstrate our theory as follows.
@@
Instead of directly calculating the magnetic moment of a single phonon, we focus on experimentally observable phonon Zeeman energy induced by the coupling between PMM and spins of electrons. We start from the total phonon Hamiltonian including the time-reversal symmetry (TRS)-breaking perturbation as
@@
where $a,b=x,y,z$ represent the Cartesian coordinates, l=1,2,Nl=1,2,...N denotes the lattice site for a system containing $N$ atoms, and $u_{l,a}=\sqrt{m}x_{l,a}$ represents the renormalized atomic displacement with $m$ being the atomic mass. The last term in Eq. (1), denoted as $\delta H$, describes the TRS-breaking perturbation resulting from spin-phonon interaction [1], and $\eta_{l,ab}$ is a $3\times 3$ real antisymmetric matrix satisfying $\eta_{l,ab}=-\eta_{l,ba}$. In addition, $p_{l,a}=\dot{u}_{l,a}+\sum_{l,b}\eta_{l,ab}u_{l,b}$ represents the canonical momentum [58]. By identifying the phonon angular momentum as $s^{\text{ph}}_{l,c}=\sum_{ab}\epsilon_{abc}u_{l,a}p_{l,b}$ (c is the direction of angular momentum), with $\epsilon_{abc}$ being the third-order Levi-Civita tensor, we can express $\delta H$ as phonon Zeeman energy in a compact form $E_{\textrm{Z}}=\bm{\eta}\cdot\bm{s}^{\text{ph}}=\sum_{l,ab}\eta_{l,c}s^{\text{ph}}_{l,c}$, where we have defined $\eta_{l,c}=\frac{1}{2}\sum_{ab}\epsilon_{abc}\eta_{l,ab}$. Notably, we find that phonon angular momentum can be treated as an effective perturbative field to the electron-phonon coupled system, and thus phonon Zeeman energy can be effectively regarded as a linear response.
Importantly, we will show that $\bm{\eta}$ can be derived from the TRS-breaking self-energy correction of phonons. According to $E_{\textrm{Z}}=\bm{\eta}\cdot\bm{s}^{\text{ph}}$, we have
$\eta_{l,c}=\frac{\partial H}{\partial s^{\text{ph}}_{l,c}}=\frac{1}{p_{l,b}}\frac{\partial H}{\partial u_{l,a}}$ (2)
where we have applied the chain derivative rule. Applying Kubo formula, we can get
$\left<\eta_{l,c}\right>=-\frac{i}{\hbar}\int_{-\infty}^{t}dt^{\prime}u_{l,a}p_{l,b}\left\langle\left[\frac{-1}{p_{l,a}}\frac{\partial H}{\partial u_{l,b}}(t),\frac{1}{p_{l,b}}\frac{\partial H}{\partial u_{l,a}}(t^{\prime})\right]\right\rangle_{0}.$ (3)
By taking Fourier transform of Eq. (3) and considering $p_{l,a}=-i\omega u_{l,a}$, we can obtain $\eta^{c}(\bm{q},\omega)$ ($\bm{q}$ is the phonon wave vector, and $\omega$ denotes phonon frequency) as
$\eta_{c}(\bm{q},\omega)=\frac{1}{\hbar\omega}\int dte^{-i\omega t}\theta(t)\left\langle\left[\hat{V}_{a}(\bm{q},\omega),\hat{V}_{b}(-\bm{q},\omega)\right]\right\rangle_{0}.$ (4)
Here, $\theta(t)$ is the Heaviside step function, and the EPC interaction operator $\hat{V}_{a}(\bm{q},\omega)=\sum_{l_{c}}z_{l_{c}}(\bm{q},\omega)\frac{\partial H}{\partial u_{l,a}}$, where $z_{l_{c}}(\bm{q},\omega)$ is the normalized contribution factor of the $l_{c}$-th atom for the phonon mode $(\bm{q},\omega)$[58].
Now, we take electron’s spin degree of freedom into consideration. Taking the spin-quantization axis of electron as the c direction, we could get the relation that $\bm{\eta}(\bm{q},\omega)=\frac{1}{i\omega}\bm{\chi}^{-}(\bm{q},\omega)$, with $\chi_{c}^{-}(\bm{q},\omega)=\chi_{c,\uparrow}(\bm{q},\omega)-\chi_{c,\downarrow}(\bm{q},\omega)$ (see SM [58] for details), indicating that the spin polarization of electrons leads to phonon Zeeman splittings. The susceptibility $\chi_{c,s}(\bm{q},\omega)$ for $s=\uparrow,\downarrow$ can be obtained by performing the Matsubara sum,
@@
Here, $\bm{k}$ represents the wave vector of the electron, $N_{\bm{k}}$ is the total number of $\bm{k}$, and $\epsilon_{s}^{\bm{k},\alpha}$ denotes the electronic energy with band indices $\alpha$. $f$ is the Fermi-Dirac distribution, and $V_{c,s}^{\bm{k},\alpha\beta}(\bm{q},\omega)$ is the EPC tensor [59, 60]. It should be emphasized that all parameters in Eq. (5) can be directly obtained from ab initio calculations, without empirical inputs or additional assumptions. A non-zero $\bm{\chi}^{-}$ requires spin-polarized electronic states, for example, in ferromagnets. To realize large $\bm{\chi}^{-}$, there are two key points according to Eq. (5): (i) The contribution of EPC tensor $V_{c,s}^{\bm{k},\alpha\beta}(\bm{q},\omega)$ is substantial, and materials composed of light atoms tend to possess strong $V_{c,s}^{\bm{k},\alpha\beta}(\bm{q},\omega)$. (ii) When the energy difference $(\epsilon^{\bm{k}+\bm{q},\alpha}-\epsilon^{\bm{k},\beta})$ approaches $\hbar\omega$, it leads to an enhancement of $\bm{\chi}^{-}$. Consequently, we can conclude that EPC-induced large phonon Zeeman splittings are most likely to occur in ferromagnetic narrow-gap semiconductors, semimetals, and metals containing light elements.
@@
Here, $\mu_{B}$ denotes the Bohr magneton of electron, $B_{c}$ is the effective magnetic field, and $g^{\text{ph}}_{c}=g^{e}\frac{\hbar}{2i\omega}\frac{\partial\chi_{c}^{+}}{\partial E}|_{E=E_{F}}$ represents the effective phonon $g$-factor (see SM [58]), where $\chi_{c}^{+}(\bm{q},\omega)=\chi_{c,\uparrow}(\bm{q},\omega)+\chi_{c,\downarrow}(\bm{q},\omega)$ with $g^{e}$ being the electronic $g$-factor involved in EPC. Equation (6) enables ab initio calculations of PMM. The dimensionless parameter $\frac{\hbar}{2i\omega}\frac{\partial\chi_{c}^{+}}{\partial E}|_{E=E_{F}}$ acts as an amplification factor dominated by the EPC tensor $V_{c,s}^{\bm{k},\alpha\beta}$ and density of electron states near the Fermi energy [Eq. (S28) in SM [58]], and it allows the $g$-factor of the circularly polarized phonon to reach the order of $g^{e}$, leading to a large PMM. This enhancement of PMM originates from electron loop currents induced by the EPC-mediated angular momentum transfer from ions to electrons, where electron loop currents are significantly enhanced owing to the tiny electron mass. In contrast, within the classical framework, the phonon’s $g$-factor is constrained to $10^{-5}\sim 10^{-4}$ order of $g^{e}$ due to large ionic masses compared to the electron mass [16, 61].
Based on the above theory, we develop a dedicated computational module within Quantum ESPRESSO [62, 63] incorporating density-functional perturbation theory with Hubbard corrections (DFPT+U) [59, 60] for precise determination of response coefficients $\bm{\eta}$ and Landé $g$-factors of phonons. Then we add the effect of $\bm{\eta}$ into Schrödinger-like phonon equations constructed from the dynamical matrix through near-degenerate perturbation theory (see SM [58] for details), yielding TRS-breaking phonon dispersions.
Comparison with experiment. A recent experimental work reported the first observation of magnetic order induced circularly polarized phonons and their spontaneous splitting in the ferromagnetic Weyl semimetal $\mathrm{Co_{3}Sn_{2}S_{2}}$ [46]. As shown in Fig. 2(a), $\mathrm{Co_{3}Sn_{2}S_{2}}$ belongs to centrosymmetric space group $\mathrm{R\bar{3}m}$, and exhibits out-of-plane ferromagnetic ordering below the Curie temperature ($\sim$175 K). Using helicity-resolved magneto-Raman scattering spectroscopy, a discernible splitting of the $E_{g}$ phonon modes at the $\Gamma$ point was observed in the ferromagnetic state, reaching up to $1.27~\mathrm{cm}^{-1}$($\sim$0.038 THz) at low temperatures (2 K). The two split modes correspond to phonon modes with opposite angular momenta, describing the right-handed and left-handed in-plane circular motion of S atoms, respectively [see Fig. 2(c) insets]. Although Ref. [46] provides an implicit EPC-based explanation of such a notable phonon Zeeman splitting, a direct and more convincing quantitative description is absent due to the lack of ab initio methods for magnetic phonon spectra.
Now we show that our ab initio method could well reproduce the experimental result quantitatively. Giving full consideration to EPC effects, for the first time, we obtain the complete magnetic phonon dispersions of $\mathrm{Co_{3}Sn_{2}S_{2}}$ throughout the Brillouin zone (see the SM for more details [58]). Figure  2(b) presents the experimentally relevant magnetic phonon dispersion along the high-symmetry path $\Gamma$-T. Remarkably, our calculations clearly demonstrate a splitting of 0.042 THz ($\mu_{\text{ph}}=2.8\times 10^{-3}\mu_{B}$) at the $\Gamma$ point for the $E_{g}$ mode [Fig. 2(c)], in good agreement with the observed value of 0.038 THz. Beyond this, we surprisingly find that the two circularly polarized phonon modes cross each other on the $\Gamma$-T path, forming a phonon Weyl point [Fig. 2(c)]. This magnetic phonon Weyl point [71] is protected by the $C_{3v}$ point group symmetry along $\Gamma$-T, which is previously inaccessible without obtaining the full magnetic phonon dispersions.
Figure 2: Lattice structure and calculated magnetic phonon dispersions of $\mathrm{Co_{3}Sn_{2}S_{2}}$. (a) The conventional cell (black lines) and primitive cell (green lines) of $\mathrm{Co_{3}Sn_{2}S_{2}}$. (b) The magnetic phonon dispersion with phonon Zeeman splittings along $\Gamma$-T path. (c) The magnetic phonon dispersion of $E_{g}$ modes. The color gradient from red to blue represents the phonon angular momentum. Upper insets: The phonon magnetic moment for the two branches of $E_{g}$ mode along $\Gamma$-T path.
Lower insets: Schematic of the two phonon modes with opposite angular momenta, corresponding to right-handed and left-handed circular motions of S atoms, respectively.
@@
Intrinsic phonon Chern states and the inertially decoupled model. The accurate calculation of phonon Zeeman splitting enables the determination of magnetic phonon spectra throughout the Brillouin zone, which is indispensable for exploring magnetic phonon topology. Following the analogy to the Chern state in the Haldane model [72], we adopt the two-dimensional (2D) honeycomb lattice [45, 73, 74] as an illustrative example. When either inversion symmetry ($I$) or TRS ($T$) is broken, each phonon Dirac cone at KK($K^{\prime}$) valley acquires a finite gap, effectively gaining masses denoted as $\Delta_{I}$ and $\Delta_{T}$, respectively. It has been shown that when $|\Delta_{T}|>|\Delta_{I}|$, band inversion occurs in one of the two Dirac cones, resulting in a topological phonon state with a Chern number of $|C|=1$ [45].
Figure 3: Inertially decoupled model (IDM) and intrinsic phonon Chern states. (a) Schematic of circularly polarized phonon modes in the high frequency regime: When $m\sim M$, both atoms contribute to high-frequency modes (upper panel); in the IDM condition ($m\ll M$), high-frequency modes are primarily contributed by light atoms (lower panel). (b) Phonon dispersions of nonmagnetic honeycomb lattice and ferromagnetic triangular lattice. (c-e) TRS-breaking phonon dispersions of the 2D IDM by combining the honeycomb and triangular lattices (c,d) and the 3D IDM by stacking 2D IDM along the $z$-direction (e). (f) Edge states (cyan/magenta) of the 2D IDM. Parameters are detailed in the SM [58].
@@
However, it is challenging to realize such phonon Chern states in realistic materials with honeycomb lattices. Although Coriolis forces and strong magnetic fields can produce a nonzero $\Delta_{T}$ [45], these approaches either require large external fields or complex experimental measurements. In contrast, the phonon Zeeman splitting could offer a sizable $\Delta_{T}$ in magnets without the need of external fields. To further satisfy the condition of $|\Delta_{T}|>|\Delta_{I}|$, it is necessary to minimize $|\Delta_{I}|$. Therefore, an elemental ferromagnetic honeycomb material inherently preserving inversion symmetry with $|\Delta_{I}|=0$ would be the ideal candidate. Unfortunately, to our knowledge, no elemental ferromagnetic honeycomb material has been discovered.
@@
To seek realistic materials that exhibit intrinsic phonon Chern states, we propose a practical composite lattice model, referred to as the inertially decoupled model that comprises two types of atoms with a significant mass (inertia) difference. A prototypical example of this model is constructed by combining a honeycomb lattice composed of nonmagnetic light atoms of mass $m$ with a triangular lattice of heavy magnetic atoms of mass $M$ ($m\ll M$), as shown in Figs. 3(b,c). Due to the difference in atomic inertia, the high-frequency and low-frequency dispersions are mainly contributed by the light and heavy atoms, respectively, as shown in Figs. 3(a,c), which is termed “inertially decoupled” (see SM [58] for details). Within the honeycomb sublattice, the inversion symmetry is preserved, resulting in a vanishing $\Delta_{I}$ mass at both the $K$ and $K^{\prime}$ valleys. Meanwhile, the magnetic $M$ atoms break TRS to generate large phonon Zeeman splittings, producing the desired large $\Delta_{T}$ mass. The large $\Delta_{T}$ opens a topologically nontrivial phonon gap (seen in Fig. 3(d)), thereby realizing an intrinsic phonon Chern state with $|C|=1$, characterized by topologically protected edge states crossing the phonon gap, as shown in Fig. 3(f).